# Disorder-dependent dominant superconducting pairing symmetry in doped graphene

### v1

@@ IV Conclusion
In summary, we utilized the DQMC method to investigate the response of the superconducting order parameter to bond disorder in doped graphene. We focused on computing the pairing susceptibility $P_{\alpha}$ and effective pairing interaction $P_{\alpha}-\widetilde{P}_{\alpha}$ for three different electron densities: $\left\langle n\right\rangle=0.95$, $\left\langle n\right\rangle=0.85$, and $\left\langle n\right\rangle=0.70$. We observed that when the electron density was near half-filling ($\left\langle n\right\rangle=0.95$), both the $d+id$ wave and $ES$ wave effective pairing correlations were positive, indicating the coexistence of both pairing waves. In addition, as the disorder strength increase, the system gradually transform from $d+id$ wave dominance to $ES$ wave dominance, suggesting that a certain strength of disorder can alter the dominant superconducting pairing. Similar behavior was observed at lower electron densities, with the distinction that for lower electron densities, both the $d+id$ and $ES$ wave effective pairing correlations were negative, indicating the absence of these two pairing symmetries. Our calculations may provide insights into the understanding of superconductivity in systems where doping and disorder coexist, as encountered in experiments.
@@ Appendix A Convergence of the mean values of measurements
The introduction of bond disorder into the system brings a certain deviation to the results due to the randomness of disorder. To ensure the accuracy of the results, it is necessary to take the average of multiple groups of disordered results. In Fig.8(a), we show the effective pairing interaction $P_{\alpha}-\widetilde{P}_{\alpha}$ with the number of disorder realizations. For any given density $\left\langle n\right\rangle$, the data does not exhibit significant changes after the number of realizations exceeds 10. We also show the variation as a function of the number of disordered realizations in Fig.8(b), and the variance curve shows good convergence. In the inset of Fig.8(b), we calculate the average of several sets of data with the number of realizations $N$ set to 5, 10, 20, and 30, respectively. The results show that the fluctuations are more severe when $N$=5, and are significantly suppressed as $N$ is further increased. This confirms the rationality of using 20 realizations in our main text.
@@ Appendix C Sign problem
Away from half-filling usually leads to sign problems. Related researches have shown that the presence of bond disorder may alleviate the sign problemPaiva et al. (2015), which is beneficial for our calculations. In Fig.10, we plot the average fermion sign $\left\langle sign\right\rangle$, which is the ratio of the integral of the product of up and down spin determinants to the integral of the absolute value of the productIglovikov et al. (2015)
@@
as a function of temperature. When $\left\langle n\right\rangle=0.95$, the system is near half-filling, and the impact of the sign problem is weak, so we mainly show the average sign for (a) $\left\langle n\right\rangle=0.85$ and (b) $\left\langle n\right\rangle=0.70$. As is seen, the sign problem is particularly serious at low temperatures. Fortunately, disorder shall limit and weaken the impact of the sign problem. In order to obtain the same quality of data as $\left\langle sign\right\rangle\approx 1$, much longer runs are necessary to compensate the fluctuations. Indeed, we can estimate that the runs need to be stretchedSantos (2003); Blankenbecler et al. (1981); Yang et al. (2016) by a factor on the order of $\left\langle sign\right\rangle^{-2}$. In our simulations, especially in the simulation results where the sign problem is much worse, we have increased measurement from 10,000 to 100,000 times to compensate the fluctuations, and thus the results for current parameters are reliable.
@@ Appendix D Finite-size effect
To make our diagram more convincing, we also checked results on different lattice size. In fig.11, we take $\left\langle n\right\rangle=0.85$, $\Delta=1.5$ and $\Delta=2.0$ as examples, and plot $P_{d+id}-\widetilde{P}_{d+id}$ as a function of temperature for different lattice sizes $L=3$, $L=4$, and $L=5$. It can be seen that the finite-size has little effect on the results, and there is no qualitative difference in the results at low temperatures, indicating that the $L=4$ we have chosen can accurately reflect the dominant pairing.

### v2

@@ IV Conclusion
In summary, we utilized the DQMC method to investigate the response of the superconducting order parameter to bond disorder in doped graphene. We focused on computing the pairing susceptibility $P_{\alpha}$ and effective pairing interaction $P_{\alpha}-\widetilde{P}_{\alpha}$ for three different electron densities: $\left\langle n\right\rangle=0.95$, $\left\langle n\right\rangle=0.85$, and $\left\langle n\right\rangle=0.70$. We observed that when the electron density was near half-filling ($\left\langle n\right\rangle=0.95$), both the $d+id$ wave and $ES$ wave effective pairing correlations were positive, indicating the coexistence of both pairing waves. In addition, as the disorder strength increases, the system gradually transforms from $d+id$ wave dominance to $ES$ wave dominance, suggesting that a certain strength of disorder can alter the dominant superconducting pairing. Similar behavior was observed at lower electron densities, with the distinction that for lower electron densities, both the $d+id$ and $ES$ wave effective pairing correlations were negative, indicating the absence of these two pairing symmetries. Our calculations may provide insights into the understanding of superconductivity in systems where doping and disorder coexist, as encountered in experiments.
@@ Appendix A Convergence of the mean values of measurements
The introduction of bond disorder into the system brings a certain deviation to the results due to the randomness of disorder. To ensure the accuracy of the results, it is necessary to take the average of multiple groups of disordered results. In Fig. 8(a), we show the effective pairing interaction $P_{\alpha}-\widetilde{P}_{\alpha}$ with the number of disorder realizations. For any given density $\left\langle n\right\rangle$, the data does not exhibit significant changes after the number of realizations exceeds ten. We also show the variation as a function of the number of disordered realizations in Fig. 8(b), and the variance curve shows good convergence. In the inset of Fig.8(b), we calculate the average of several sets of data with the number of realizations $N$ set to 5, 10, 20, and 30, respectively. The results show that the fluctuations are more severe when $N$=5, and are significantly suppressed as $N$ is further increased. This confirms the rationality of using 20 realizations in our main text.
@@ Appendix C Sign problem
Away from half-filling usually leads to sign problems. Related researches have shown that the presence of bond disorder may alleviate the sign problem [55], which is beneficial for our calculations. In Fig.10, we plot the average fermion sign $\left\langle sign\right\rangle$, which is the ratio of the integral of the product of up and down spin determinants to the integral of the absolute value of the product [56]
@@
as a function of temperature. When $\left\langle n\right\rangle=0.95$, the system is near half-filling, and the impact of the sign problem is weak, so we mainly show the average sign for (a) $\left\langle n\right\rangle=0.85$ and (b) $\left\langle n\right\rangle=0.70$. As is seen, the sign problem is particularly serious at low temperatures. Fortunately, disorder shall limit and weaken the impact of the sign problem. In order to obtain the same quality of data as $\left\langle sign\right\rangle\approx 1$, much longer runs are necessary to compensate the fluctuations. Indeed, we can estimate that the runs need to be stretched [57, 58, 59] by a factor on the order of $\left\langle sign\right\rangle^{-2}$. In our simulations, especially in the simulation results where the sign problem is much worse, we have increased measurement from 10000 to 100000 times to compensate the fluctuations, and thus, the results for current parameters are reliable.
@@ Appendix D Finite-size effect
To make our diagram more convincing, we also checked results on different lattice size. In Fig.11, we take $\left\langle n\right\rangle=0.85$, $\Delta=1.5$, and $\Delta=2.0$ as examples, and plot $P_{d+id}-\widetilde{P}_{d+id}$ as a function of temperature for different lattice sizes $L=3$, $L=4$, and $L=5$. It can be seen that the finite size has little effect on the results, and there is no qualitative difference in the results at low temperatures, indicating that the $L=4$ we have chosen can accurately reflect the dominant pairing.